OpenStudy (aleah54):

simplification of the expression (xyz^2)^4. Can you explain how you got your answer

OpenStudy (mathmate):

hint: Use the law of exponents: \((a\times b \times c)^n\) \(=a^n\times b^n \times c^n\)